A 10.000-in-diameter plunger slides in a 10.006-in-diameter cylinder, the annular space being filled with oil having a kinematic viscosity of 0.004 ft2/s and specific gravity of 0.85. If the plunger moves at 0.6ft/s, find the force, in lbf, when 9 ft is engaged in the cylinder


1
Expert's answer
2021-03-10T05:53:06-0500

"Plunger \space diameter=10.000 in"

"Cylinder \space diameter=10.006 in"

"Kinematic \space viscosity =0.004ft^2/s"

"G=0.85"

Plunger moves 0.6ft/s

"\rho=\mu(\frac{du}{dx}), \delta=\frac{0.85 \times62.4}{32.2}=1.647 plunges/ft^3"

"\mu=1.647 \times0.004=0.006588 lb.s/ft^2"

"du=\frac{10.006-10.000}{2}/12=0.000250 ft"

"Z=0.006588(0.6\/0.00025)=15.81 lb\/ft^2"

"F_f=ZA=15.81\times(9 \times \pi(\frac{10}{12})=372.5 lb"